40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 920     Entry time: Thu Sep 4 07:46:10 2008
Author: Yoichi 
Type: Update 
Category: IOO 
Subject: MC is now happy 
The MC has been locked for more than 12 hours continuously now !
Changes I made yesterday were:
(1) Removed the 20dB attenuator before the EOM.
(2) Reduced the Fast Gain from 21dB to 16dB, which made the PC to be a little bit more loaded (~0.6Vrms).

As Rana pointed out in the meeting, setting the Fast Gain a bit lower may have put the FSS in a stabler state.
Attachment 1: MC-lock.png  63 kB  | Hide | Hide all
MC-lock.png
ELOG V3.1.3-